Main page‎ > ‎Observing Runs‎ > ‎Night Reports‎ > ‎

Night report page - 20130207


Observers: Nancy Elias-Rosa, Antonia Morales-Garoffolo
Data reduction team: Sandra Benitez, Stefan Taubenberger

Observations with SOFI.
Cloudy at South at the beginning of the night. Humidity between 57% and 70% all night.

Log

**The electricity in La Silla fell down again (yesterday the same thing happened). Despite having a generator, the passage from one to another has made the TO to restart all the telescopes and instruments. 

SOFI calibration data done in the afternoon

**The guide camera has some problem that the TO could not resolve this afternoon. So, we are going to observe just with the tracking of the telescope. Therefore, we cannot do image analysis (we need the guide for this). After the electricity cut, we do not know how EFOSC2 can work if we shift to this instrument because this has not being tested in the afternoon. 

--DIMM SEEING 0.97"-- 

UT 00:25      Focus 

UT 00:35      Hip010502              airmass=1.32        BG       Telluric STD for SN2012ec.

UT 00:50      SN2012ec               airmass=1.36        BG

--DIMM SEEING 0.96"-- 

UT 02:20      Hip036187              airmass=1.61        BG RG    Telluric STD for OGLE-2012-SN-006.

UT 02:40      OGLE-2012-SN-006       airmass=1.65        BG       The SN is too faint.

UT 05:35      Hip014112              airmass=2.31        BG       Telluric STD for OGLE-2012-SN-006.

--DIMM SEEING 0.58"--

UT 05:40      SN2013K                airmass=2.17        JHKs

UT 06:27      GD153                  airmass=1.91        BG       Flux STD. OB template with low exposures. 

--DIMM SEEING 0.98"--

UT 06:38      PSNJ15213475-0722183   airmass=2.14        JHKs    

--DIMM SEEING 1.02"--

UT 07:51      EG274                  airmass=1.55        BG      Flux STD.
  
UT 07:58      LSQ13fn                airmass=1.01        JHKs

UT 09:20      END of the night.

Good night



Comments